Solve the system by addition. 30) x - 2y = -16 -7x - ly = -8
Perform the indicated operations. 31) ( 2x2 + 2x - x3 + 1) - (4x2 - 8x - x3 + 6)
Factor the difference of squares. 32) 49x2 - 16
Perform the indicated operation and simplify. ܕ܂ .3 ܙ30
Write a slope-intercept equation for a line with the given characteristics. 34) Passes through (3,-7) and (0,-2)
Answer #1

5019 30 24-2y = -10 - - 74-19 =-8 - multiply eyn o by 7 8 adding with eqn 6 weget 12-14y= -112 + -724-1y = -8 -15y = -120 - y